# OpenOmni: A Collaborative Open Source Tool for Building Future-Ready Multimodal Conversational Agents

### v1

@@ Abstract
Multimodal conversational agents are highly desirable because they offer natural and human-like interaction. However, there is a lack of comprehensive end-to-end solutions to support collaborative development and benchmarking. While proprietary systems like GPT-4o and Gemini demonstrating impressive integration of audio, video, and text with response times of 200-250ms, challenges remain in balancing latency, accuracy, cost, and data privacy. To better understand and quantify these issues, we developed OpenOmni, an open-source, end-to-end pipeline benchmarking tool that integrates advanced technologies such as Speech-to-Text, Emotion Detection, Retrieval Augmented Generation, Large Language Models, along with the ability to integrate customized models. OpenOmni supports local and cloud deployment, ensuring data privacy and supporting latency and accuracy benchmarking. This flexible framework allows researchers to customize the pipeline, focusing on real bottlenecks and facilitating rapid proof-of-concept development. OpenOmni can significantly enhance applications like indoor assistance for visually impaired individuals, advancing human-computer interaction. Our demonstration video is available https://www.youtube.com/watch?v=zaSiT3clWqY, demo is available via https://openomni.ai4wa.com, code is available via https://github.com/AI4WA/OpenOmniFramework.
@@ 1 Introduction
The ideal form of multimodal HCI should mirror human interactions, incorporating video and audio inputs with audio outputs. Despite the availability of various modular components, there is no comprehensive integrated open-source implementation to foster research and innovation in this field. Integrating existing models—such as audio speech recognition (Speech2Text), multimodal large models (MLMs), and text-to-speech synthesis (TTS)—into a multimodal conversation system reveals significant challenges in balancing latency and accuracy. Historically, accuracy has been a major hurdle; however, advancements in large language models (LLMs) have substantially improved contextual relevance. The main challenge is reducing end-to-end latency while maintaining accuracy. While OpenAI and Google have shown it’s possible, the open-source community lacks alternatives that replicate this performance.
@@
To support the rapid and responsible development of this new HCI format, establishing robust evaluation and benchmarking protocols is essential. For instance, if a user initiates a conversation in a sad and urgent tone, the system should respond appropriately with patient. Evaluating this interaction is crucial and challenging for widespread adoption. Our project aims to bridge these gaps by:
@@ 2 Related works
Recently, Kyutai, a technology company, from France released a planned open-source, fully end-to-end multimodal conversational AI called Moshi  [^5] https://kyutai.org/. This model supports text and audio modalities, excluding images, and claims to achieve an end-to-end latency of 200ms. We can integrate the video modality via an Image2Text Lin et al. (2021) module into Moshi, creating a Hybrid solution, as shown in Figure 4, that combines the divide-and-conquer and fully end-to-end approaches. Another feasible Hybrid solution is to use speech-to-text to convert audio into text, then feed this text along with video (processed into image sequences) to a vision language model, which generates text responses. These responses can then be processed through text-to-speech, as illustrated in Figure 2 via the Vision LLM line.
@@ 4.2 Can “AI” be your president?
One intensive conversational scenario is a debate. We extracted segments from the US Presidential Debate 2024 between Biden and Trump [^8] https://www.youtube.com/watch?v=-v-8wJkmwBY, focusing on Biden addressing the public and handling questions. These segments were fed into our pipeline to evaluate its performance under different configurations: OpenAI Whisper for speech-to-text, GPT-4o vision model, and text-to-speech (GPT4O_ETE); a locally deployed quantization LLM with Whisper, text-to-speech, and our emotion detection model for video input (QuantizationLLM_ETE); a version using HuggingFace LLM for inference (HF_ETE); and a version using only Whisper, GPT-3.5, and text-to-speech, ignoring the video modality (GPT35_ETE). We ran the Agent modules on an NVIDIA-3080 GPU with 12GB memory.
@@
After annotation with our interface, accuracy statistics are automatically generated. As shown in Figure 7, the average score for each conversation is 2.4. Text-to-speech can be improved with more natural emotion or personality. The generated content is often too general and sometimes inappropriate. Biden’s responses are more in-context and evidence-supported. The pipeline excelled only in answering a subjective question about Biden’s age, where the GPT-4o pipeline performed well. The GPT35_ETE pipeline had the best overall accuracy, but its responses were often in-context yet pompous. Thus, Biden still outperforms AI. In conclusion, “AI cannot be the President of the US just yet, considering both latency and accuracy.”☺
@@ 4.3 Assist the visually impaired
While latency and the need for external information currently preventing AI from mission critical tasks, conversational agents can be production-ready and useful for non-latency-critical areas that do not require extensive external knowledge. Assisting indoor activities for the visually impaired is one such application. We prepared questions for the visually impaired, including locating objects, navigating indoors, and inquiries about the surroundings. Six questions were sampled and fed to the GPT4O_ETE pipeline. One scenario demonstration is included in our provided YouTube video. The latency statistics in Figure 8 show responses within approximately 30 seconds.

### v2

@@ Abstract
Multimodal conversational agents are highly desirable because they offer natural and human-like interaction. However, there is a lack of comprehensive end-to-end solutions to support collaborative development and benchmarking. While proprietary systems like GPT-4o and Gemini have demonstrated impressive integration of audio, video, and text with response times of 200-250ms, challenges remain in balancing latency, accuracy, cost, and data privacy. To better understand and quantify these issues, we developed OpenOmni, an open-source, end-to-end pipeline benchmarking tool that integrates advanced technologies such as Speech-to-Text, Emotion Detection, Retrieval Augmented Generation, Large Language Models, along with the ability to integrate customized models. OpenOmni supports local and cloud deployment, ensuring data privacy and supporting latency and accuracy benchmarking. This flexible framework allows researchers to customize the pipeline, focusing on real bottlenecks and facilitating rapid proof-of-concept development. OpenOmni can significantly enhance applications like indoor assistance for visually impaired individuals, advancing human-computer interaction. Our demonstration video is available https://www.youtube.com/watch?v=zaSiT3clWqY, demo is available via https://openomni.ai4wa.com, code is available via https://github.com/AI4WA/OpenOmniFramework.
@@ 1 Introduction
The ideal form of multimodal HCI should mirror human interactions, incorporating video and audio inputs with audio outputs. Despite the availability of various modular components, there is no comprehensive integrated open-source implementation to foster research and innovation in this field. Integrating existing models, such as audio speech recognition (Speech2Text), multimodal large models (MLMs), and text-to-speech synthesis (TTS)—into a multimodal conversation system reveals significant challenges in balancing latency and accuracy. Historically, accuracy has been a major hurdle; however, advancements in large language models (LLMs) have substantially improved contextual relevance. The main challenge is reducing end-to-end latency while maintaining accuracy. While OpenAI and Google have shown it’s possible, the open-source community lacks alternatives that replicate this performance.
@@
To support the rapid and responsible development of this new HCI format, establishing robust evaluation and benchmarking protocols is essential. For instance, if a user initiates a conversation in a sad and urgent tone, the system should respond appropriately with patience. Evaluating this interaction is crucial and challenging for widespread adoption. Our project aims to bridge these gaps by:
@@ 2 Related works
Recently, Kyutai, a technology company from France, released a planned open-source, fully end-to-end multimodal conversational AI called Moshi  [^5] https://kyutai.org/. This model supports text and audio modalities, excluding images, and claims to achieve an end-to-end latency of 200ms. We can integrate the video modality via an Image2Text Lin et al. (2021) module into Moshi, creating a Hybrid solution, as shown in Figure 4, that combines the divide-and-conquer and fully end-to-end approaches. Another feasible Hybrid solution is to use speech-to-text to convert audio into text, then feed this text along with video (processed into image sequences) to a vision language model, which generates text responses. These responses can then be processed through text-to-speech, as illustrated in Figure 2 via the Vision LLM line.
@@ 4.2 Can “AI” be your president?
One intensive conversational scenario is a debate. We extracted segments from the US Presidential Debate 2024 between Biden and Trump [^8] https://www.youtube.com/watch?v=-v-8wJkmwBY, focusing on Biden addressing the public and handling questions. After downloading the videos, you can use a prepared script in our codebase to split them into segments. This script allows you to specify the start and end times of each conversation, enabling you to create a conversational dataset from the videos. These segments were fed into our pipeline to evaluate its performance under different configurations: OpenAI Whisper for speech-to-text, GPT-4o vision model, and text-to-speech (GPT4O_ETE); a locally deployed quantization LLM with Whisper, text-to-speech, and our emotion detection model for video input (QuantizationLLM_ETE); a version using HuggingFace LLM for inference (HF_ETE); and a version using only Whisper, GPT-3.5, and text-to-speech, ignoring the video modality (GPT35_ETE). We ran the Agent modules on an NVIDIA-3080 GPU with 12GB memory.
@@
After annotation with our interface, accuracy statistics are automatically generated. The accuracy metrics here include evaluation metrics like WER, CER Roy (2021) for speech2text task, overall scores given by the annotators, etc. As shown in Figure 7, the average score for each conversation is 2.4. Text-to-speech can be improved with more natural emotion or personality. The generated content is often too general and sometimes inappropriate. Biden’s responses are more in-context and evidence-supported. The pipeline excelled only in answering a subjective question about Biden’s age, where the GPT-4o pipeline performed well. The GPT35_ETE pipeline had the best overall accuracy, but its responses were often in-context yet pompous. Thus, Biden still outperforms AI. In conclusion, “AI cannot be the President of the US just yet, considering both latency and accuracy.”☺
@@ 4.3 Assist the visually impaired
While latency and the need for external information currently preventing AI from mission critical tasks, conversational agents can be production-ready and useful for non-latency-critical areas that do not require extensive external knowledge. Assisting indoor activities for the visually impaired is one such application, in which you can either utilize high-speed internet or limit data transfer to local exchanges. These type applications can benefit from maintaining high input/output rates, helping to mitigate latency issues. We prepared questions for the visually impaired, including locating objects, navigating indoors, and inquiries about the surroundings. Six questions were sampled and fed to the GPT4O_ETE pipeline. One scenario demonstration is included in our provided YouTube video. In this scenario, video and audio data stream from the client side and are saved to storage along with exportable metadata accessible via the admin portal. This setup allows you to export annotated datasets, including raw video and audio data, for developing new models. The latency statistics in Figure 8 show responses within approximately 30 seconds.